# Hierarchical Working Memory and a New Magic Number

### v1

@@ Abstract
The extremely limited working memory span, typically around four items, contrasts sharply with our everyday experience of processing much larger streams of sensory information concurrently. This disparity suggests that working memory can organize information into compact representations such as chunks, yet the underlying neural mechanisms remain largely unknown. Here, we propose a recurrent neural network model for chunking within the framework of the synaptic theory of working memory. We showed that by selectively suppressing groups of stimuli, the network can maintain and retrieve the stimuli in chunks, hence exceeding the basic capacity. Moreover, we show that our model can dynamically construct hierarchical representations within working memory through hierarchical chunking. A consequence of this proposed mechanism is a new limit on the number of items that can be stored and subsequently retrieved from working memory, depending only on the basic working memory capacity when chunking is not invoked. Predictions from our model were confirmed by analyzing single-unit responses in epileptic patients and memory experiments with verbal material. Our work provides a novel conceptual and analytical framework for understanding the on-the-fly organization of information in the brain that is crucial for cognition.
@@ Introduction
Working memory (WM) is hypothesized to be a distinct capacity for holding and manipulating multiple pieces of information, which is crucial for human cognitive abilities such as verbal communication, reading comprehension, and abstract reasoning [1, 2, 3]. Paradoxically, however, people typically cannot simultaneously hold more than four items in WM [4]. For example, repeating several words or digits is practically effortless and mistake-free, but for lists of five random words, people begin making mistakes [5, 6, 7]. How, then, are people able to process much larger streams of inputs, such as long passages of text or movies? One attractive idea is chunking, i.e., organizing several items into higher-level units [8, 9, 10, 11, 12, 13]. Sometimes chunks are stored in long-term memory due to previous experience [14, 15], e.g., when observing a stream of letters, a combination like “CIA” can be processed as a separate unit rather than simply a group of three letters. Another example is a familiar expression like “Oh my God” or “Easier said than done”. These pre-existing chunks could be thought of as having stable memory representations and could therefore be encoded and processed as a single item. Conceptually more challenging is the phenomenon of spontaneous chunking, where novel combinations of items are grouped into separate units “on the fly”, as when a phone number is divided into chunks of 2-3 digits each, or words in a sentence are combined into units based on their syntactic role, such as “–a little boy–was dressed – in a green shirt”. Indeed, this sentence is much easier to remember than a random sequence of nine words. Surprisingly, a minor manipulation like introducing slight pauses between presentations of consecutive groups of items is enough to trigger chunking and the corresponding increase in capacity [16, 17, 18, 19]. In this study, we addressed two inter-related questions inspired by the above considerations: how spontaneous chunking might emerge in the brain and what is (if there is one) the limit for the number of items that can be held in WM when spontaneous chunking is activated.
@@ Results
The RNN detailed in Methods Sec. .1 exhibits different dynamical regimes, depending on the STP parameters and external background input. In particular, as shown in [24, 26], there exists a regime of cyclic behavior where items that were loaded into the network via external stimuli are maintained in WM in the form of sequential brief reactivations called population spikes [33]. As the number of loaded memories increases, the network eventually fails to maintain some of them, i.e., there is a maximal number of items that can be maintained in the WM, $C$, which depends on the synaptic-level parameters of the RNN [26].
@@
(b) Simulation of the network in (a). $R^{(k)}$: activity trace of firing rates, color-coded to match the corresponding clusters in (a). The time-course of the traces is labeled as chunks (stimulus clusters), pauses (chunking clusters), and long pauses (meta-chunking clusters). $I^{b(k)}$: activity traces of background input currents. Decreasing the background input to a cluster at level $k$ suppresses its reactivation and removes the inhibition on its children clusters at level $k-1$.
@@
The main idea of the proposed chunking mechanism is that at no point in time do more than a small number of stimulus clusters exhibit activity reactivations in the form of population spikes, thus not exceeding the basic WM capacity. The crucial point is that due to synaptic augmentation, stimulus clusters that are currently suppressed still have stronger recurrent self-connections than the ones that were not active at a given trial as long as augmentation has not disappeared. Therefore, the network can retrieve temporarily suppressed items by consecutively switching off the chunking clusters, releasing the suppressed stimulus clusters within a corresponding chunk from inhibition. To illustrate this mechanism, we simulate a network of 16 clusters (both stimulus and chunking), 6 of which are activated consecutively with transient external input (presentation stage, the shades in Fig. 1(d)).
@@
We first consider continuous presentation of 6 inputs to the stimulus clusters with no chunking activated. At the end of the presentation, 4 of the corresponding clusters remain active in the form of periodic population spikes while two other clusters drop out of WM, corresponding to a WM capacity of 4 for the chosen values of parameters, similar to [26] (the top panel of Fig. 1(d)). Now consider presenting the same six memory items, but with a slightly longer interval between the presentation of the $3^{rd}$ and $4^{th}$ items, during which a chunking cluster is activated (shown in red in the bottom panel of Fig. 1(d)). We assume that the chunking cluster quickly inhibits the three stimulus clusters that were presented before it (the three blue colors) and remains the only cluster active until the items of the next chunk are presented to the network (shown in green). A second chunking cluster is then activated, shown in purple. This way, the network effectively binds stimulus clusters in each chunk with its corresponding chunking cluster (Fig. 1(b)). We assume that the fast inhibition between chunking clusters and corresponding stimulus clusters happens through strengthening the existing dense but weak inhibitory synapses between them (Fig. 1(a)).
@@
After all stimuli are presented, the network maintains reactivations of two chunking clusters while the synaptic variables of the stimulus clusters slowly decay to their baseline values. However, if the chunking cluster is suppressed within the augmentation time-window $\tau_{A}$, the items that were inhibited by it will bounce back (Fig. 1(d) bottom panel, the blue colors in the retrieval stage). At this point in time, four clusters are active: the second chunking cluster and three stimulus clusters from the first chunk, with all items from the first chunk being successfully retrieved. When the second chunk is to be retrieved, the first chunking cluster is again activated by control input while the second chunking cluster is suppressed, allowing the stimulus clusters from the second chunk to activate. This chunking scenario allows the retrieval of all six memory items while at any given moment in time, the network maintains no more than four active clusters, not exceeding basic WM capacity. In this way, chunking increases effective working memory capacity by reducing the concurrent load on working memory, at the expense of activating higher-level representations (chunking clusters). Note that although here we model the chunking cues as slight pauses in-between the presentations, in general, chunking could be triggered by other cues, such as tonic variations and semantic meanings. The idea that chunking reduces the load on working memory was first introduced in the psychology literature [14, 37, 9]. Subsequently, neuroimaging studies observed that chunking reduces neural activity in upstream brain regions that process raw stimuli but increases activity in downstream regions associated with higher-level representations [38, 39, 18], which is consistent with our proposed mechanism.
@@
(b) Average firing rates of non-boundary neurons over all trials for individual subjects. The subjects is sorted based on the location of the dip. A similar trend as in (a) is observed for each individual subject. For individual 2D plots, also see Fig. S3.
@@
Thanks to advances in single-neuron recording technologies, we can now test our hypothesis using data collected from drug-resistant epilepsy patients [40]. Consider the experiment reported in [40], where subjects are asked to watch a series of movie clips, each consisting of two episodes separated by a “cut” in the middle of the movie. Such movie cuts serve to induce cognitive boundaries for event segmentation in episodic memory. The authors in [40] identified a group of neurons in the medial temporal lobe that fire selectively at these boundaries and termed them “cognitive boundary” neurons. If these neurons segment episodic memories in a manner similar to how chunking clusters segment working memory in our model, then we should also observe a decrease in the firing rates of the stimulus neurons upon the firing of the cognitive boundary neurons. In [40], although the boundary neurons can be unambiguously identified by aligning their response to the movie cuts, it is difficult to pinpoint stimulus neurons due to the continuous nature of the visual stimulus. Therefore, we study the putative effect of boundary neurons on the rest of the system by aggregating neurons that are detected but not classified as boundary neurons. Upon averaging over subjects and trials, we find that after the peak in the firing rate of boundary neurons (top panel of Fig. 3 (a)), about $\sim 130$ ms later, there is a dip in the average activities of the rest of the recorded neurons (bottom panel of Fig. 3 (a)). Furthermore, there is a continuous ramp-up of activities following the dip. This trend can also be seen at the level of individual subjects (Fig. 3 (b)), and qualitatively agrees with the prediction of our hierarchical working memory model.
@@
An important prediction of the hierarchical working memory model is the existence of an absolute limit $M^{*}$, beyond which perfect retrieval is impossible (Eq. (1)). One of the earliest studies to quantify this transition is the experiment performed by Miller and Selfridge [41] on the statistical approximation of language. In this experiment, the authors constructed $n$-gram approximations to English, where $n$ refers to coherent occurrences with the previous $n-1$ words. For example, a $1$-gram approximation would consist of words randomly chosen from a corpus. In a $2$-gram approximation, each word would appear coherently with the previous word, but coherence for any sliding window of three words is not required. As $n$ increases, the constructed text gradually approaches natural text. In [41], subjects were presented with verbal materials constructed from such $n$-gram approximations and asked to recall the words. The fraction of recalled words $f$ decreases with the length $L$ of the material and increases with the degree of approximation $n$ (Fig. 4(a) inset). Here, we are interested in the critical length $L_{c}$ beyond which retrieval begins to be imperfect, i.e., $f(L_{c})=1$. Since the defining feature of working memory is the ability to perfectly retrieve items that are sustained in the memory, $L_{c}$ is a measure of working memory capacity.
@@ Acknowledgements.
Acknowledgments. We want to thank Tankut Can, Antonis Georgiou, Máté Lengyel, Gianluigi Mongillo, Sandro Romani, Sebastian Seung, and Jie Zheng for helpful discussions. W.Z. is supported by the Starr Foundation at the Institute for Advanced Study. M.T. is supported by the Israeli Science Foundation grant 1657/19 and Foundation Adelis.
@@ Methods
Activity traces of all the dynamical variables in Eq. (3)-(8) are shown in the Fig. S1. In particular, the synaptic matrix $J_{\mu\nu}$ before and after chunking in Fig. 1(d) is shown for comparison. All simulation parameters are reported in Table 1. All the external inputs $I^{e}$ used for loading the memories are rectangular functions with support only at the presentation time, and have an amplitude of $750\ \mathrm{Hz}$, and all the background input $I^{b}$ has amplitude of $|I^{b}|=10\ \mathrm{Hz}$. Additionally, the timing of the external control signals are summarized in below.
Fig. 1(d) top panel: Stimulus starts to load at $t=1\ \mathrm{s}$ for a duration of $0.025\ \mathrm{s}$ with an interval of $0.45\ \mathrm{s}$. Background input $I^{b}$ has a constant value of $10\ \mathrm{Hz}$.
Fig. 1(d) bottom panel: Stimulus starts to load at $t=1\ \mathrm{s}$ for a duration of $0.025\ \mathrm{s}$ with an interval of $0.45\ \mathrm{s}$. Chunking clusters are loaded for a duration of $0.025\ \mathrm{s}$ with an interval of $0.3\ \mathrm{s}$. Background input $I^{b}$ has a constant value of $10\ \mathrm{Hz}$ during the presentation stage and switches between $10\ \mathrm{Hz}$ and $-10\ \mathrm{Hz}$ for a duration of $1.35\ \mathrm{s}$ during the retrieval stage.
Fig. 2(b): The $k=3$ level stimulus clusters start to load at $t=1\ \mathrm{s}$ for a duration of $0.15\ \mathrm{s}$ with an interval of $0.45\ \mathrm{s}$. $k=2,3$ level chunking clusters load for a duration of $0.01\ \mathrm{s}$ with an interval of $0.2\ \mathrm{s}$. Background input $I^{b}$ has a constant value of $10\ \mathrm{Hz}$ during the presentation stage and switches between $10\ \mathrm{Hz}$ and $-10\ \mathrm{Hz}$ for a duration of $0.8\ \mathrm{s}$ during the retrieval stage.
@@
     Eq. (6)-(7) $\tau$ Membrane time constant $8\ \mathrm{ms}$
$w^{EI}$ Synaptic efficacy $I\to E$ 1.5
$w^{IE}$ Synaptic efficacy $E\to I$ 2.4
$\alpha$ Gain function parameter $1.5\ \mathrm{Hz}$
$P$ Total number of clusters 16
     Eq. (3)-(5) $\tau_{f}$ Facilitation time constant $1.2\ \mathrm{s}$
$\tau_{d}$ Depression time constant $0.45\ \mathrm{s}$
$\tau_{A}$ Augmentation time constant $75\ \mathrm{s}$
$U$ Facilitation baseline $0.3$
$A_{\text{min}}$ Augmentation baseline 8
$A_{\text{max}}$ Maximum augmentation 30
$\kappa_{A}$ Augmentation rate 0.03
     Eq. (8) $J_{\text{inh}}$ Inhibition strength $10$
@@
(b) Snapshots of the synaptic matrix $J_{\mu\nu}$ before and after chunking. Clusters 1-14 are stimulus clusters and 15-16 are chunking clusters. At $t=2\ \mathrm{s}$, only the first chunk (the blue colors) is presented, chunking clusters are not activated, and only the recurrent self-connections are nonzero in the synaptic matrix. At $t=4.5\ \mathrm{s}$, both chunks are formed by the chunking connections (dark blue colors in the top right corner).
@@
Individual subjects’ z-score firing rates of the non-boundary neurons are shown in blue, with one standard deviation included as shades. Black dashed lines denote $t=0\ \mathrm{s}$ where the movie cut occurs. Red dashed lines denote the location of the maximum firing rate of the boundary neurons. Results are pooled from the raw firing rates of all non-boundary neurons from that subject. Subject IDs are presented according to the data in [40]. While some subjects do not exhibit the qualitative trend as predicted (e.g., the firing rate of subject P64CS does not have a ramp, and TWH120 does not have a dip), most of the subjects’ firing rates follow the same qualitative trend as observed in the average plot in Fig. 3(a).

### v2

# Synaptic Theory of Chunking in Working Memory

@@ Abstract
Working memory often appears to exceed its basic span by organizing items into compact representations called chunks. Chunking can be learned over time for familiar inputs; however, it can also arise spontaneously for novel stimuli. Such on-the-fly structuring is crucial for cognition, yet the underlying neural mechanism remains unclear. Here we introduce a synaptic theory of chunking, in which short-term synaptic plasticity enables the formation of chunk representations in working memory. We show that a specialized population of “chunking neurons” selectively controls groups of stimulus-responsive neurons, akin to gating. As a result, the network maintains and retrieves the stimuli in chunks, thereby exceeding the basic capacity. Moreover, we show that our model can dynamically construct hierarchical representations within working memory through hierarchical chunking. A consequence of this proposed mechanism is a new limit on the number of items that can be stored and subsequently retrieved from working memory, depending only on the basic working memory capacity when chunking is not invoked. Predictions from our model were confirmed by analyzing single-unit responses in epileptic patients and memory experiments with verbal material. Our work provides a novel conceptual and analytical framework for understanding how the brain organizes information in real time.
@@ Introduction
Working memory (WM) is hypothesized to be a distinct capacity for holding and manipulating multiple pieces of information, which is crucial for human cognitive abilities such as verbal communication, reading comprehension, and abstract reasoning [1, 2, 3]. Paradoxically, however, people typically cannot simultaneously hold more than four items in WM [4]. For example, repeating several words or digits is practically effortless and mistake-free, but for lists of five random words, people begin making mistakes [5, 6, 7]. How, then, are people able to process much larger streams of inputs, such as long passages of text or movies? One attractive idea is chunking, i.e., organizing several items into higher-level units [8, 9, 10, 11, 12, 13]. Sometimes chunks are stored in long-term memory due to previous experience [14, 15], e.g., familiar expression like “Oh my God” or “Easier said than done” can be processed as a coherent unit rather than individual words. These pre-existing chunks could be thought of as having stable memory representations learned and consolidated over time, and could therefore be encoded and processed as a single item. However, conceptually more challenging is the phenomenon of spontaneous chunking, where novel combinations of items are grouped into separate units “on the fly”, as when a phone number is divided into chunks of 2-3 digits each, or words in a sentence are combined into units based on their syntactic role, such as “–a little boy–was dressed – in a green shirt”. Indeed, this sentence is much easier to remember than a random sequence of nine words. Surprisingly, a minor manipulation like introducing slight pauses between presentations of consecutive groups of items is enough to trigger chunking and the corresponding increase in capacity [16, 17, 18, 19]. In this study, we addressed two inter-related questions inspired by the above considerations: how spontaneous chunking might emerge in the brain and what is (if there is one) the limit for the number of items that can be held in WM when spontaneous chunking is activated.
@@ Results
The RNN detailed in Methods Sec. .1 exhibits different dynamical regimes, depending on the STP parameters and external background input. In particular, as shown in [24, 26], at high background input level, there exists a persistent activity regime where clusters have sustained elevated firing rates corresponding to loaded memory items. As the background input is lowered, there exists a low-activity regime with cyclic behavior where items that were loaded into the network via external stimuli are maintained in WM in the form of sequential brief reactivations called population spikes [33]. As the number of loaded memories increases, the network eventually fails to maintain some of them, i.e., there is a maximal number of items that can be maintained in the WM, $C$, which depends on the synaptic-level parameters of the RNN [26].
@@
(b) Simulation of the network in (a). $R^{(k)}$: activity trace of firing rates, color-coded to match the corresponding clusters in (a). The time-course of the traces is labeled as chunks (stimulus clusters), pauses (chunking clusters), and long pauses (meta-chunking clusters). $I^{b(k)}$: activity traces of background input currents. Decreasing the background input to a cluster at level $k$ suppresses its reactivation and removes the inhibition on its children clusters at level k1k-1.
@@
The main idea of the proposed chunking mechanism is that the chunking clusters can selectively activate and suppress the stimulus clusters, so that at no point in time do more than a small number of stimulus clusters reactivate as population spikes, thus not exceeding the basic WM capacity. Due to synaptic augmentation, stimulus clusters that are currently suppressed by the chunking clusters still have stronger recurrent self-connections than the ones that were not active at a given trial as long as augmentation has not disappeared. Therefore, the network can retrieve temporarily suppressed items by sequentially switching off the chunking clusters, releasing the suppressed stimulus clusters within the corresponding chunk from inhibition.
@@
To demonstrate the chunking mechanism, we simulate a network of 16 clusters (both stimulus and chunking), 6 of which are activated consecutively with transient external input (presentation stage, the shades in Fig. 1(d)). We first consider continuous presentation of 6 inputs to the stimulus clusters with no chunking activated. At the end of the presentation, 4 of the corresponding clusters remain active in the form of periodic population spikes while two other clusters drop out of WM, corresponding to a WM capacity of 4 for the chosen values of parameters, similar to [26] (the top panel of Fig. 1(d)). Now consider presenting the same six memory items, but with a slightly longer interval between the presentation of the $3^{rd}$ and $4^{th}$ items, during which a chunking cluster is activated (shown in red in the bottom panel of Fig. 1(d)). We assume that the chunking cluster quickly inhibits the three stimulus clusters that were presented before it (the three blue colors) and remains the only cluster active until the items of the next chunk are presented to the network (shown in green). A second chunking cluster is then activated, shown in purple. This way, the network effectively binds the stimulus clusters in each chunk to their corresponding chunking cluster (Fig. 1(b)). Such group-specific binding is akin to gating [37], where the activity of each chunking cluster gates the entire chunk of stimulus clusters via inhibition.
@@
We assume that the fast inhibition between chunking clusters and corresponding stimulus clusters happens through strengthening the existing dense but weak inhibitory synapses between them (Fig. 1(a)). After all stimuli are presented, the network maintains reactivations of two chunking clusters while the synaptic variables of the stimulus clusters slowly decay to their baseline values. However, if the chunking cluster is suppressed within the augmentation time-window $\tau_{A}$, the items that were inhibited by it will bounce back (Fig. 1(d) bottom panel, the blue colors in the retrieval stage). At this point in time, four clusters are active: the second chunking cluster and three stimulus clusters from the first chunk, with all items from the first chunk being successfully retrieved. When the second chunk is to be retrieved, the first chunking cluster is again activated by control input while the second chunking cluster is suppressed, allowing the stimulus clusters from the second chunk to activate. This chunking scenario allows the retrieval of all six memory items while at any given moment in time, the network maintains no more than four active clusters, not exceeding basic WM capacity. In this way, chunking increases effective working memory capacity by reducing the concurrent load on working memory, at the expense of activating higher-level representations (chunking clusters).
Above, we chose to illustrate the chunking mechanism in the periodic activity regime because the mechanistic effects of chunking clusters are most apparent with regular firing traces. Nevertheless, our proposed chunking mechanism applies to both the persistent-activity and periodic-activity regimes, with chunking clusters serving the same function in each. Note that, although we model the chunking cues here as slight pauses between presentations, in general chunking can be triggered by other cues, such as tonic variations and semantic meanings. The idea that chunking reduces the load on working memory was first introduced in the psychology literature [14, 38, 9]. Subsequently, neuroimaging studies observed that chunking reduces neural activity in upstream brain regions that process raw stimuli but increases activity in downstream regions associated with higher-level representations [39, 40, 18], which is consistent with our proposed mechanism.
@@
(b) Average firing rates of non-boundary neurons over all trials for individual subjects. Subjects are sorted based on the location of the dip. A trend similar to panel (a) is observed for each subject. For individual 2D plots, see Fig. S3.
(c) Average firing rates of neurons aligned to the onset of the movie (relative time zero). After the peak in onset-specific neurons, the non-onset-specific neurons do not exhibit the dip-then-ramp pattern seen in panel (a). Top: onset-specific neurons. Bottom: Non-onset specific neurons.
@@
Thanks to advances in single-neuron recording technologies, we can now test our hypothesis using data collected from drug-resistant epilepsy patients [41]. Consider the experiment reported in [41], where subjects are asked to watch a series of movie clips, each consisting of two episodes separated by a “cut” in the middle of the movie. Such movie cuts serve to induce cognitive boundaries for event segmentation in episodic memory. The authors in [41] identified a group of neurons in the medial temporal lobe that fire selectively at these boundaries and termed them “cognitive boundary” neurons. If these neurons segment episodic memories in a manner similar to how chunking clusters segment working memory in our model, then we should also observe a decrease in the firing rates of the stimulus neurons upon the firing of the cognitive boundary neurons. In [41], although the boundary neurons can be unambiguously identified by aligning their responses to movie cuts, it is difficult to pinpoint stimulus neurons due to the continuous nature of the visual stimulus. Therefore, we study the putative effect of boundary neurons on the rest of the system by aggregating neurons that are detected but not classified as boundary neurons. We align all the neurons to the movie cuts. Upon averaging over subjects and trials, we find that after the peak in the firing rate of boundary neurons (top panel of Fig. 3 (a)), about $\sim 130$ ms later, there is a dip in the average activities of the rest of the recorded neurons (bottom panel of Fig. 3 (a)). Furthermore, there is a continuous ramp-up of activities following the dip. This trend is also evident at the level of individual subjects (Fig. 3 (b)), and qualitatively agrees with the prediction of our hierarchical working memory model. As a control, within the same recorded population we label the subset of neurons that respond to the onset of the movie clip as “onset” neurons. Aligning firing rates to the movie onset, we observe a peak in the onset neurons as reported in [41]; however, unlike Fig. 3 (a), the remaining neurons (including boundary neurons) do not exhibit the dip-then-ramp pattern (Fig. 3 (c)). This indicates that the dip and ramp-up are specific to boundary neurons and suggests an internal network mechanism rather than simple inhibitory feedback or statistical artifacts.
@@
An important prediction of the hierarchical working memory model is the existence of an absolute limit $M^{*}$, beyond which perfect retrieval is impossible (Eq. (1)). One of the earliest studies to quantify this transition is the experiment performed by Miller and Selfridge [42] on the statistical approximation of language. In this experiment, the authors constructed $n$-gram approximations to English, where $n$ refers to coherent occurrences with the previous n1n-1 words. For example, a $1$-gram approximation would consist of words randomly chosen from a corpus. In a $2$-gram approximation, each word would appear coherently with the previous word, but coherence for any sliding window of three words is not required. As $n$ increases, the constructed text gradually approaches natural text. In [42], subjects were presented with verbal materials constructed from such $n$-gram approximations and asked to recall the words. The fraction of recalled words $f$ decreases with the length $L$ of the material and increases with the degree of approximation $n$ (Fig. 4(a) inset). Here, we are interested in the critical length $L_{c}$ beyond which retrieval begins to be imperfect, i.e., $f(L_{c})=1$. Since the defining feature of working memory is the ability to perfectly retrieve items that are sustained in the memory, $L_{c}$ is a measure of working memory capacity.
@@ Acknowledgements.
Acknowledgments. All authors were involved in the conceptualization of the study. WZ conducted the numerical simulations and data analysis. All authors contributed to writing the manuscript. The authors declare no competing interests. All data analyzed in the manuscript are publicly available online. The simulation code will be made available in a GitHub repository upon publication.
We want to thank Tankut Can, Antonis Georgiou, Máté Lengyel, Gianluigi Mongillo, Sandro Romani, Sebastian Seung, and Jie Zheng for helpful discussions. W.Z. is supported by Eric and Wendy Schmidt Member in Biology Fund and the Simons Foundation at the Institute for Advanced Study. MK is supported in part by a grant from Fran Morris Rosman and Richard Rosman. M.T. is supported by the Simons Foundation, MBZUAI-WIS Joint Program for Artificial Intelligence Research and Foundation Adelis.
@@ Methods
Activity traces of all the dynamical variables in Eq. (3)-(8) are shown in the Fig. S1. In particular, the synaptic matrix $J_{\mu\nu}$ before and after chunking in Fig. 1(d) is shown for comparison. All simulation parameters are reported in Table 1. All the external inputs $I^{e}$ used for loading the memories are rectangular functions with support only at the presentation time, and have an amplitude of 750Hz750\ $\mathrm${Hz}, and all the background input $I^{b}$ has amplitude of $|I^{b}|=10\ \mathrm{Hz}$. Additionally, the timing of the external control signals are summarized in below.
Fig. 1(d) top panel: Stimulus starts to load at t=1st=1\ $\mathrm${s} for a duration of 0.025s0.025\ $\mathrm${s} with an interval of 0.45s0.45\ $\mathrm${s}. Background input $I^{b}$ has a constant value of 10Hz10\ $\mathrm${Hz}.
Fig. 1(d) bottom panel: Stimulus starts to load at t=1st=1\ $\mathrm${s} for a duration of 0.025s0.025\ $\mathrm${s} with an interval of 0.45s0.45\ $\mathrm${s}. Chunking clusters are loaded for a duration of 0.025s0.025\ $\mathrm${s} with an interval of 0.3s0.3\ $\mathrm${s}. Background input $I^{b}$ has a constant value of 10Hz10\ $\mathrm${Hz} during the presentation stage and switches between 10Hz10\ $\mathrm${Hz} and 10Hz-10\ $\mathrm${Hz} for a duration of 1.35s1.35\ $\mathrm${s} during the retrieval stage.
Fig. 2(b): The $k=3$ level stimulus clusters start to load at t=1st=1\ $\mathrm${s} for a duration of 0.15s0.15\ $\mathrm${s} with an interval of 0.45s0.45\ $\mathrm${s}. $k=2,3$ level chunking clusters load for a duration of 0.01s0.01\ $\mathrm${s} with an interval of 0.2s0.2\ $\mathrm${s}. Background input $I^{b}$ has a constant value of 10Hz10\ $\mathrm${Hz} during the presentation stage and switches between 10Hz10\ $\mathrm${Hz} and 10Hz-10\ $\mathrm${Hz} for a duration of 0.8s0.8\ $\mathrm${s} during the retrieval stage.
@@
   Eq. (6)-(7) $\tau$ - Membrane time constant 8ms8\ $\mathrm${ms}
$w^{EI}$ - Synaptic efficacy $I\to E$ 1.5
$w^{IE}$ - Synaptic efficacy $E\to I$ 2.4
$\alpha$ - Gain function parameter 1.5Hz1.5\ $\mathrm${Hz}
$P$ - Total number of clusters 16
   Eq. (3)-(5) $\tau_{f}$ - Facilitation time constant 1.2s1.2\ $\mathrm${s}
$\tau_{d}$ - Depression time constant 0.45s0.45\ $\mathrm${s}
$\tau_{A}$ - Augmentation time constant 75s75\ $\mathrm${s}
$U$ - Facilitation baseline $0.3$
$A_{\text{min}}$ - Augmentation baseline 8
$A_{\text{max}}$ - Maximum augmentation 30
$\kappa_{A}$ - Augmentation rate 0.03
   Eq. (8) $J_{\text{inh}}$ - Inhibition strength 1010
@@
(b) Snapshots of the synaptic matrix $J_{\mu\nu}$ before and after chunking. Clusters 1-14 are stimulus clusters and 15-16 are chunking clusters. At t=2st=2\ $\mathrm${s}, only the first chunk (the blue colors) is presented, chunking clusters are not activated, and only the recurrent self-connections are nonzero in the synaptic matrix. At t=4.5st=4.5\ $\mathrm${s}, both chunks are formed by the chunking connections (dark blue colors in the top right corner).
@@
Individual subjects’ z-score firing rates of the non-boundary neurons are shown in blue, with one standard deviation included as shades. Black dashed lines denote t=0st=0\ $\mathrm${s} where the movie cut occurs. Red dashed lines denote the location of the maximum firing rate of the boundary neurons. Results are pooled from the raw firing rates of all non-boundary neurons from that subject. Subject IDs are presented according to the data in [41]. While some subjects do not exhibit the qualitative trend as predicted (e.g., the firing rate of subject P64CS does not have a ramp, and TWH120 does not have a dip), most of the subjects’ firing rates follow the same qualitative trend as observed in the average plot in Fig. 3(a).